# Large Vision-Language Model Alignment and Misalignment: A Survey Through the Lens of Explainability

### v2

@@ A Survey Through the Lens of Explainability
Dong Shu1, Haiyan Zhao2, Jingyu Hu3, Weiru Liu3, Ali Payani4, Lu Cheng5, Mengnan Du2 1Northwestern University, 2New Jersey Institute of Technology, 3University of Bristol, 4Cisco Research,5University of Illinois Chicago dongshu2024@u.northwestern.edu,  mengnan.du@njit.edu
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have achieved significant progress in multimodal understanding. Models such as GPT-4V OpenAI , Gemini Team et al. (2023), LLaVA Liu et al. (2024b), Claude-3.5-Sonnet Anthropic (2024), Qwen2-VL Wang et al. (2024d), and LLaMa 3.2 Dubey et al. (2024) demonstrate remarkable capabilities in tasks like image captioning and visual question answering, not only processing visual and textual information independently but also reasoning across these modalities. These advances are built upon two two key foundations: large language models (LLMs) and vision encoders. LLMs such as GPT-3.5 Brown (2020), LLaMA Touvron et al. (2023a), LLaMA 2 Touvron et al. (2023b), and Qwen Bai et al. (2023a) transformed natural language processing, while visual understanding models like Vision Transformer (ViT) Dosovitskiy et al. (2021) and vision-language models like CLIP Radford et al. (2021) have advanced the ability to process visual information and create aligned visual-textual representations respectively.
@@
In this survey, we present a structured framework for understanding and addressing alignment challenges in LVLMs from an explainability perspective. We first examine the fundamentals of alignment, including its representational and behavioral aspects, training procedures, and theoretical foundations. We then analyze misalignment phenomena across three semantic levels: object, attribute, and relational misalignment. Our investigation reveals that misalignment stems from challenges at the data level (e.g., quality and balance issues), model level (e.g., architectural limitations and ability gaps), and inference level (e.g., task discrepancies). We review existing mitigation strategies and outline future directions, emphasizing the need for standardized evaluation protocols and in-depth explainability studies.
@@ 2 Alignment of LVLMs
In this section, we examine alignment in LVLMs across four essential dimensions. First, we define the concept of alignment in LVLMs. Second, we detail the procedural stages through which alignment is achieved in practice. Third, we explore the theoretical foundations that make alignment possible between visual and textual modalities. Finally, we discuss methods for measuring and evaluating alignment in LVLMs in Appendix B.
@@ 2.1 What is Alignment?
2.2 How is Alignment Achieved?
The development of alignment in LVLMs progresses through three major stages (see Figure 3 in Appendix), each is built upon its predecessor to achieve increasingly sophisticated cross-modal integration. Additionally, we present prominent real-world LVLM architectures in Appendix A.
Stage 2: Adapter Fine-tuning.  The second stage involves fine-tuning an adapter module that bridges the pre-trained visual encoder with the language model. This stage introduces lightweight adapter architectures, which typically consist of simple components such as linear layers, MLPs, or cross-attention layers that learn to translate between visual and language model embedding spaces. For example, cross-attention layers can feed image encoder representations into the language model, enabling the model to attend to relevant visual features when generating text Meta AI (2024). A key characteristic of this approach is the preservation of the original capabilities of both the visual encoder and language model while learning to interface between them. During adapter training, while the visual encoder parameters may be updated, the language model parameters often remain frozen to maintain their original text capabilities. This intermediate stage is essential for establishing effective connections between modalities while preserving the specialized capabilities of each component.
Stage 3: End-to-End Fine-tuning.  The final stage involves comprehensive fine-tuning of the entire system, including the visual encoder, adapter, and LLM components together. This comprehensive approach allows for deeper integration and more sophisticated alignment between all components. It enables the model to learn task-specific optimizations that require coordinated adjustments across all modules. Through this process, the model develops more advanced cross-modal understanding capabilities and facilitates the emergence of emergent behaviors that arise from the deep integration of visual and textual processing. This stage often results in the highest performance but requires careful balancing to avoid catastrophic forgetting or degradation of pre-existing capabilities.
@@ 2.3 Why is Alignment Possible?
Theoretical Perspective.  From a theoretical standpoint, visual and textual data are different projections of the same underlying reality. As Huh et al. argue in their Platonic Representation Hypothesis Huh et al. (2024), all modalities are measurements of a real world that generates our observations. When humans create images or write text, they are encoding information about this same reality, though through different measurement processes. Although these modalities appear distinct on the surface, they fundamentally capture overlapping semantic information about the same world state. This shared origin in physical reality, combined with the fact that humans generate both types of data to describe their observations of the world, provides the theoretical foundation for why these modalities can be meaningfully aligned in a common representation space.
@@ 3.1 Definition of Misalignment
Misalignment in LVLMs occurs when the model’s output semantically diverges from the visual content it is meant to describe. These discrepancies show in several key phenomena, impacting the overall performance of these models. In this paper, we categorize behavior-level misalignment phenomena in LVLM into three semantic levels $\mathcal{S}=\{s_{o},s_{a},s_{r}\}$: object misalignment ($s_{o}$), attribute misalignment ($s_{a}$), and relation misalignment ($s_{r}$). Rather than using the term ‘hallucination’ commonly found in the literature Liu et al. (2024a), we adopt the term ‘misalignment’ to better characterize how these discrepancies emerge between visual and language representations.
@@
To better illustrate real-life examples of misalignment, we have selected four prominent LVLMs and provided their generated responses in Appendix D.
@@ 3.2.1 Dataset Level
Data quality and distribution patterns play crucial roles in contributing to misalignment between visual and language representations in LVLMs. Several key dataset factors can impede the model’s ability to form accurate associations between visual inputs and textual descriptions, affecting both training effectiveness and inference performance. We list one factor below, for additional factors please refer to Appendix C.
@@ 3.2.2 Model Level
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14.7em,font=,, where level=2text width=8.0em,font=,, where level=3text width=9.5em,font=,, where level=4text width=12em,font=,, [Mitigation Methods (§4),ver [Parameter-Tuning Alignment (§4.1),ver [Improving Training
Scheme [Contrastive
Learning, text width=6em [e.g., CIT Hu et al. (2023), HACL Jiang et al. (2024) , leaf, text width=42em] ] [Instruction
Tuning, text width=6em [e.g., IDK-Instructions Cha et al. (2024), REVERIE Zhang et al. (2025), TextSquare
 Tang et al. (2024), LRV-Instruction Liu et al. (2023), AIT Park et al. (2024), MMINSTRUCT
 Liu et al. (2024d) , leaf, text width=42em] ] [RLHF, text width=6em [e.g., Fact-RLHF Sun et al. (2023), RLHF-V Yu et al. (2024c) , leaf, text width=42em] ] [Preference
Optimization, text width=6em [e.g., CLIP-DPO Ouali et al. (2025), HA-DPO Zhao et al. (2023b), FDPO Gunjal et al. (2024),
HSA-DPO Xiao et al. (2024), HalluciDoctor Yu et al. (2024b), Chen et al. (2023b), ReCaption
 Wang et al. (2024c), MOCHa Ben-Kish et al. (2024), SILKIE Li et al. (2023a), V-DPO
 Xie et al. (2024), RoVRM Wang et al. (2024a) , leaf, text width=42em] ] ] [Improving Model
Architecture [Vision
Encoder, text width=6em [e.g., InternVL Chen et al. (2024c), Ferret You et al. (2023), Vcoder Jain et al. (2024), Monkey
 Li et al. (2024b), ObjMLM Dai et al. (2022) , leaf, text width=42em] ] [Connection
Module, text width=6em [e.g., PATCH Shang et al. (2024), HallE-Switch Zhai et al. (2023) , leaf, text width=42em] ] ] ] [Parameter-Frozen Alignment (§4.2),ver [Augment-Based [Retrieval, text width=6em [e.g., ARA Qu et al. (2024), RAM Chen et al. (2023a), LMCap Ramos et al. (2023a), Smallcap
 Ramos et al. (2023b), Sarto et al. (2024), Pensieve Yang et al. (2024) , leaf2, text width=42em] ] [Generate, text width=6em [e.g., MARINE Zhao et al. (2024d), Li et al. (2024a), Zhao et al. (2023a), RITUAL
 Woo et al. (2024a), Kim et al. (2024b), VDGD Ghosh et al. (2024) , leaf2, text width=42em] ] ] [Inference-Based [e.g., VTI Liu et al. (2024c), ICT Chen et al. (2024a), MetaToken Fieback et al. (2024) , leaf2, text width=49.6em] ] [Decoding-Based [e.g., AVC Woo et al. (2024b), DAMRO Gong et al. (2024), VCD Leng et al. (2024), ICD Wang et al. (2024e),
CODE Kim et al. (2024a), PAI Liu et al. , OPERA Huang et al. (2024a), SGD Min et al. (2024), Han et al. (2024),
Yue et al. (2024), IBD Zhu et al. (2024a), HALC Chen et al. (2024b), DBD Feng et al. (2024), AGLA
 An et al. (2024), RVD Zhong et al. (2024), M3ID Favero et al. (2024), DeCo Wang et al. (2024b), CGD
 Deng et al. (2024) , leaf2, text width=49.6em] ] [Post-Decoding [e.g., LURE Zhou et al. (2023b), Woodpecker Yin et al. (2023), Yu et al. (2024a), LogicCheckGPT Wu et al. (2024a),
VOLCANO Lee et al. (2023) , leaf2, text width=49.6em] ] ] ]
@@ 4 Mitigation Methods
Building upon our analysis of misalignment causes in LVLMs, we now examine strategies for mitigating these challenges (see Figure 2). These mitigation approaches can be categorized into two groups: parameter-tuning alignment methods and parameter-frozen alignment methods. Parameter-tuning alignment involves modifying specific components within the LVLM architecture to reduce misalignment through targeted parameter updates. In contrast, parameter-frozen alignment methods address misalignment while maintaining the LVLM’s original parameters unchanged, offering solutions that preserve the model’s structure while improving its cross-modal alignment capabilities. We have conducted experiments comparing the performance and computational cost of several mitigation methods (see Appendix E for details).
@@ 4.1 Parameter Tuning Alignment
Improving Training Scheme.  Parameter-tuning methods that improve the training scheme often address misalignment broadly as a data-level issue or as a general visual-textual misalignment Ouali et al. (2025); Jiang et al. (2024). This understanding leads to a straightforward objective, which is reducing the modality gap between visual and textual representations. This can often achieved by improving the dataset quality or optimizing training techniques. One common approach is contrastive learning, exemplified by methods such as CIT Hu et al. (2023) and HACL Jiang et al. (2024). These techniques involve using a third model to generate positive and negative data pairs. The LVLM is then trained to bring the representations of positive pairs closer together while pushing negative pairs apart in the embedding space. Another widely adopted strategy is instruction tuning, as seen in LRV-Instruction Liu et al. (2023) and TextSquare Tang et al. (2024). Similarly, these approaches rely on a third model to generate instructional data, which is subsequently used to train the LVLM effectively. However, these approaches often lack robust quality assurance mechanisms to verify the accuracy or relevance of the generated data, introducing potential risks. Alternatively, Reinforcement Learning from Human Feedback (RLHF) employs human feedback to train a reward model, ensuring that the generated data aligns with human preferences Sun et al. (2023); Yu et al. (2024c). While RLHF guarantees high-quality training data, it comes at a significant cost. To address this, some methods leverage preference optimization, wherein multiple responses are generated for the same input image, ranked or scored by a third model, and categorized into positive and negative pairs Ouali et al. (2025); Zhao et al. (2023b); Gunjal et al. (2024). The model is then fine-tuned on this curated dataset. Although these methods can significantly improve the model, they are often constrained by either high resource requirements (as in RLHF) or the uncertain quality of generated data (as in contrastive learning and instruction tuning) or rerank model (as in preference optimization). This highlights the ongoing need for large, diverse, and high-quality datasets to effectively address data-level misalignment.
@@ 4.2 Parameter Frozen Alignment
Augment-based Methods.  As analyzed in Section 3, insufficient input of image information is one of the primary causes of misalignment, leading to poor visual understanding. To address this, retrieval-augmented generation (RAG) methods have been adapted to dynamically integrate external knowledge into LVLMs through retrieved results Qu et al. (2024); Chen et al. (2023a); Ramos et al. (2023a, b); Sarto et al. (2024); Yang et al. (2024). By reranking the similarity of image-text pairs, RAG approaches provide more visual context and guidance to the model. Similarly, other methods rely on generating approach to enrich the input with additional information. For instance, Zhao et al. (2024d), Li et al. (2024a), and Zhao et al. (2023a) propose integrating an auxiliary model to generate relevant information based on the image. Alternatively, methods such as RITUAL Woo et al. (2024a) bypass the need for external models. It enhances the model’s exposure to diverse visual contexts by applying random transformations to input images. Additionally, approaches like Kim et al. (2024b); Ghosh et al. (2024) employ self-generated textual descriptions appended to the input prompt, ensuring the model has sufficient knowledge to answer questions accurately.

### v3

@@ A Survey Through the Lens of Explainability
Dong Shu1, Haiyan Zhao2, Jingyu Hu3, Weiru Liu3, Ali Payani4, Lu Cheng5, Mengnan Du2thanks: Corresponding author. 1Northwestern University, 2New Jersey Institute of Technology, 3University of Bristol, 4Cisco Research,5University of Illinois Chicago dongshu2024@u.northwestern.edu,  mengnan.du@njit.edu
@@ 1 Introduction
Recent Large Vision-Language Models (LVLMs) have achieved significant progress in multimodal understanding. Models such as GPT-4V OpenAI , Gemini Team et al. (2023), LLaVA Liu et al. (2024b), Claude-3.5-Sonnet Anthropic (2024), Qwen2-VL Wang et al. (2024d), and LLaMa 3.2 Dubey et al. (2024) demonstrate remarkable capabilities in tasks like image captioning and visual question answering, not only processing visual and textual information independently but also reasoning across these modalities. These advances are built upon two key foundations: large language models (LLMs) and vision encoders. LLMs such as GPT-3.5 Brown (2020), LLaMA Touvron et al. (2023a), LLaMA 2 Touvron et al. (2023b), and Qwen Bai et al. (2023a) transformed natural language processing, while visual understanding models like Vision Transformer (ViT) Dosovitskiy et al. (2021) and vision-language models like CLIP Radford et al. (2021) have advanced the ability to process visual information and create aligned visual-textual representations respectively.
@@
1.1 Contribution and Uniqueness
Our Contributions. In this survey, we present a structured framework for understanding and addressing alignment challenges in LVLMs from an explainability perspective. Our major contributions are listed as follows: (1) We examine the fundamentals of alignment, covering its representational and behavioral aspects, training procedures, and theoretical foundations (Section 2). (2) We analyze misalignment at both the representation and behavior levels, and categorize behavioral phenomena into object, attribute, and relational misalignment (Section 3). (3) We identify that misalignment arises from challenges at three key levels: data level, model level, and inference level (Section 3). (4) We review existing mitigation strategies with an emphasis on their underlying motivations, trade-offs, and assumptions, and further enhance the explainability analysis of these approaches (Section 4). (5) In the Appendix, we provide extended discussions including existing evaluation methods, real-world examples of LVLM misalignment, and mitigation methods comparative analysis.
Differences with Existing Surveys. While several existing surveys focus primarily on hallucination in LVLMs Liu et al. (2024a); Bai et al. (2024); Sahoo et al. (2024), our work addresses the broader and more foundational concept of misalignment. We view hallucination as one possible consequence of misalignment. Another key distinction is our focus on explainability as the central lens for understanding alignment. We systematically analyze alignment through the questions of what alignment is, how it is achieved, and why it is possible. This explainability-driven perspective guides our entire analysis, including the design of our mitigation taxonomy and the in-depth discussion of existing mitigation methods.
@@ 2 Alignment of LVLMs
In this section, we examine alignment in LVLMs across four essential dimensions. First, we define the concept of alignment in LVLMs. Second, we detail the procedural stages through which alignment is achieved in practice. Third, we explore the theoretical foundations that make alignment possible between visual and textual modalities. Finally, we discuss methods for measuring and evaluating alignment in LVLMs in Appendix C.
@@ 2.2 Why is Alignment Possible?
Theoretical Perspective.  From a theoretical standpoint, visual and textual data are different projections of the same underlying reality. As Huh et al. argue in their Platonic Representation Hypothesis Huh et al. (2024), all modalities are measurements of a real world that generates our observations. When humans create images or write text, they are encoding information about this same reality, though through different measurement processes. Although these modalities appear distinct on the surface, they fundamentally capture overlapping semantic information about the same world state. This shared origin in physical reality, combined with the fact that humans generate both types of data to describe their observations of the world, provides the theoretical foundation for why these modalities can be meaningfully aligned in a common representation space. See Appendix D for a formal justification.
@@ 3.1 Definition of Misalignment
Misalignment in LVLMs occurs when the model’s output semantically diverges from the visual content it is meant to describe. These discrepancies show in several key phenomena, impacting the overall performance of these models. In this section, we categorize behavior-level misalignment phenomena in LVLM into three semantic levels $\mathcal{S}=\{s_{o},s_{a},s_{r}\}$: object misalignment ($s_{o}$), attribute misalignment ($s_{a}$), and relation misalignment ($s_{r}$) (Figure 1 right). Rather than using the term ‘hallucination’ commonly found in the literature Liu et al. (2024a), we adopt the term ‘misalignment’ to better characterize how these discrepancies emerge between visual and language representations. For representation-level misalignment, please refer to Figure 1 left and Appendix B.
@@
To better illustrate real-life examples of misalignment, we have selected four prominent LVLMs and provided their generated responses in Appendix F.
@@ 3.2.1 Dataset Level
Data quality and distribution patterns play crucial roles in contributing to misalignment between visual and language representations in LVLMs. Several key dataset factors can impede the model’s ability to form accurate associations between visual inputs and textual descriptions, affecting both training effectiveness and inference performance. We list one factor below, for additional factors please refer to Appendix E.
@@ 3.2.2 Model Level
[…]
Mitigation Methods (§4)
Parameter-Tuning Alignment (§4.1) Improving Training Scheme Contrastive Learning e.g., CIT Hu et al. (2023), HACL Jiang et al. (2024) Instruction Tuning e.g., IDK-Instructions Cha et al. (2024), REVERIE Zhang et al. (2025), TextSquare Tang et al. (2024), LRV-Instruction Liu et al. (2023), AIT Park et al. (2024), MMINSTRUCT Liu et al. (2024d) RLHF e.g., Fact-RLHF Sun et al. (2023), RLHF-V Yu et al. (2024c) Preference Optimization e.g., CLIP-DPO Ouali et al. (2025), HA-DPO Zhao et al. (2023b), FDPO Gunjal et al. (2024), HSA-DPO Xiao et al. (2024), HalluciDoctor Yu et al. (2024b), Chen et al. (2023b), ReCaption Wang et al. (2024c), MOCHa Ben-Kish et al. (2024), SILKIE Li et al. (2023a), V-DPO Xie et al. (2024), RoVRM Wang et al. (2024a) Improving Model Architecture Vision Encoder e.g., InternVL Chen et al. (2024c), Ferret You et al. (2023), Vcoder Jain et al. (2024), Monkey Li et al. (2024b), ObjMLM Dai et al. (2022) Connection Module e.g., PATCH Shang et al. (2024), HallE-Switch Zhai et al. (2023) Parameter-Frozen Alignment (§4.2) Augment-Based Retrieval e.g., ARA Qu et al. (2024), RAM Chen et al. (2023a), LMCap Ramos et al. (2023a), Smallcap Ramos et al. (2023b), Sarto et al. (2024), Pensieve Yang et al. (2024a) Generate e.g., MARINE Zhao et al. (2024d), Li et al. (2024a), Zhao et al. (2023a), RITUAL Woo et al. (2024a), Kim et al. (2024b), VDGD Ghosh et al. (2024) Inference-Based e.g., VTI Liu et al. (2024c), ICT Chen et al. (2024a), MetaToken Fieback et al. (2024) Decoding-Based e.g., AVC Woo et al. (2024b), DAMRO Gong et al. (2024), VCD Leng et al. (2024), ICD Wang et al. (2024e), CODE Kim et al. (2024a), PAI Liu et al. , OPERA Huang et al. (2024a), SGD Min et al. (2024), Han et al. (2024), Yue et al. (2024), IBD Zhu et al. (2024a), HALC Chen et al. (2024b), DBD Feng et al. (2024), AGLA An et al. (2024), RVD Zhong et al. (2024), M3ID Favero et al. (2024), DeCo Wang et al. (2024b), CGD Deng et al. (2024) Post-Decoding e.g., LURE Zhou et al. (2023b), Woodpecker Yin et al. (2023), Yu et al. (2024a), LogicCheckGPT Wu et al. (2024a), VOLCANO Lee et al. (2023)
@@ 4 Mitigation Methods
Building upon our analysis of misalignment causes in LVLMs, we now examine strategies for mitigating these challenges (see Figure 2). These mitigation approaches can be categorized into two groups: parameter-tuning alignment methods and parameter-frozen alignment methods. Parameter-tuning alignment involves modifying specific components within the LVLM architecture to reduce misalignment through targeted parameter updates. In contrast, parameter-frozen alignment methods address misalignment while maintaining the LVLM’s original parameters unchanged, offering solutions that preserve the model’s structure while improving its cross-modal alignment capabilities.
@@ 4.1 Parameter Tuning Alignment
Improving Training Scheme.  Parameter-tuning methods that improve the training scheme often address misalignment broadly as a data-level issue or as a general visual-textual misalignment Ouali et al. (2025); Jiang et al. (2024). This understanding leads to a straightforward objective, which is reducing the modality gap between visual and textual representations. This can often achieved by improving the dataset quality or optimizing training techniques. One common approach is contrastive learning, exemplified by methods such as CIT Hu et al. (2023) and HACL Jiang et al. (2024). These techniques involve using a third model to generate positive and negative data pairs. The LVLM is then trained to bring the representations of positive pairs closer together while pushing negative pairs apart in the embedding space. Another widely adopted strategy is instruction tuning, as seen in LRV-Instruction Liu et al. (2023), VideoCon Bansal et al. (2024), HACA Zhao et al. (2025), and TextSquare Tang et al. (2024). Similarly, these approaches rely on a third model to generate instructional data, which is subsequently used to train the LVLM effectively. However, these approaches often lack robust quality assurance mechanisms to verify the accuracy or relevance of the generated data, introducing potential risks. Alternatively, Reinforcement Learning from Human Feedback (RLHF) employs human feedback to train a reward model, ensuring that the generated data aligns with human preferences Sun et al. (2023); Yu et al. (2024c). While RLHF guarantees high-quality training data, it comes at a significant cost. To address this, some methods leverage preference optimization, wherein multiple responses are generated for the same input image, ranked or scored by a third model, and categorized into positive and negative pairs Ouali et al. (2025); Zhao et al. (2023b); Gunjal et al. (2024). The model is then fine-tuned on this curated dataset. Although these methods can significantly improve the model, they are often constrained by either high resource requirements (as in RLHF) or the uncertain quality of generated data (as in contrastive learning and instruction tuning) or rerank model (as in preference optimization). This highlights the ongoing need for large, diverse, and high-quality datasets to effectively address data-level misalignment.
@@
POPE-Random POPE-Adversarial POPE-Popular Comp. Cost
Baseline
Mitigations
@@ 4.2 Parameter Frozen Alignment
Augment-based Methods.  As analyzed in Section 3, insufficient input of image information is one of the primary causes of misalignment, leading to poor visual understanding. To address this, retrieval-augmented generation (RAG) methods have been adapted to dynamically integrate external knowledge into LVLMs through retrieved results Qu et al. (2024); Chen et al. (2023a); Ramos et al. (2023a, b); Sarto et al. (2024); Yang et al. (2024a). By reranking the similarity of image-text pairs, RAG approaches provide more visual context and guidance to the model. Similarly, other methods rely on generating approach to enrich the input with additional information. For instance, Zhao et al. (2024d), Li et al. (2024a), and Zhao et al. (2023a) propose integrating an auxiliary model to generate relevant information based on the image. Then inject these information into the input prompt. Alternatively, methods such as RITUAL Woo et al. (2024a) bypass the need for external models. It enhances the model’s exposure to diverse visual contexts by applying random transformations to input images. Additionally, approaches like Kim et al. (2024b); Ghosh et al. (2024) employ self-generated textual descriptions appended to the input prompt, ensuring the model has sufficient knowledge to answer questions accurately.
@@
4.3 Mitigation Performance Comparison
As shown in Table 1, we selected LLaVa-V1.5-7B as the baseline model and evaluated it on POPE benchmarks. We then tested three mitigation methods: SoM-LLaVA Yan et al. (2024), SID Huo et al. (2024), and LogicCheckGPT Wu et al. (2024a), to assess their effectiveness and efficiency in mitigating misalignment in the baseline model. The baseline model was loaded directly from Hugging Face, with parameters in float16 precision and all other implementation details set to default. Each mitigation method was implemented strictly according to its official GitHub repository, without modifications to core components, except for necessary adjustments such as path configurations. More detailed analysis regarding the mitigation performance please refer to Section G in the Appendix.